Khuzawayo Construction a local sub-contractor expects to use 19 675 pockets of cement for 2021. Carrying costs are R 6.25 per unit and order costs are R 43.55. The organization works 315 Days per year. Determine the length of the order cycle.


1
Expert's answer
2021-08-31T05:43:02-0400

The number of pockets of cement that is equivalent to carrying cost=(19675*6.25)/43.55

=2823.6

=2824 full pockets of cement.

Total pockets=2824+19675

=22,499 pockets of cement

Order Cycle=(315*24*60*60)/22,499

=1,209.65secs/order